### v2

# hard problem of consciousness

Abstract
We tackle the hard problem of consciousness taking the naturally selected, self-organising, embodied organism as our starting point. We provide a mathematical formalism describing how biological systems self-organise to hierarchically interpret unlabelled sensory information according to valence and specific needs. Such interpretations imply behavioural policies which can only be differentiated from each other by the qualitative aspect of information processing. Selection pressures favour systems that can intervene in the world to achieve homeostatic and reproductive goals. Quality is a property arising in such systems to link cause to affect to motivate real world interventions. This produces a range of qualitative classifiers (interoceptive and exteroceptive) that motivate specific actions and determine priorities and preferences. Building upon the seminal distinction between access and phenomenal consciousness, our radical claim here is that phenomenal consciousness without access consciousness is likely very common, but the reverse is implausible. To put it provocatively: death grounds meaning, and Nature does not like zombies. We formally describe the multilayered architecture of self-organisation from rocks to Einstein, illustrating how our argument applies in the real world. We claim that access consciousness at the human level is impossible without the ability to hierarchically model i) the self, ii) the world/others and iii) the self as modelled by others. Phenomenal consciousness is therefore required for human-level functionality. Our proposal lays the foundations of a formal science of consciousness, deeply connected with natural selection rather than abstract thinking, closer to human fact than zombie fiction.
keywords:
hard problem of consciousness
1 Introduction
Why is anything conscious? Both biological and other physical [^1] Physical here just means non-biological. We are not suggesting biological systems are non-physical. systems process information, yet it seems that humans consciously experience in addition to merely processing information. Why? Living organisms are constantly processing self and world related information to survive in an ever-changing world. Human bodies share with all other physical systems the property of being instantiated in time and space, (e.g., our body occupies a given position and volume in space at a given time). Yet, unlike physical systems, biological, living systems are dissipative systems using energy to self-organise in the face of entropic decay and environmental perturbation [nagel1974, chalmers1995].
Initially formally introduced in the field of cybernetics [ashby1947, foerster1960] the notion of self-organisation has been applied to various disciplines including physics [haken1983], biology [camazine2003, bell2007] and neuroscience [kelso1997, Friston2010, tognoli2014]. Self-organization is typically defined as the spontaneous emergence of spatiotemporal order or pattern-formation processes in physical and biological systems resulting from interactions of its components with the environment [camazine2001, seeley2002, rosas2018].
Interestingly, much self and world related information processing goes on behind the scenes, or “in the dark” so to speak, that is, without being constantly present to our conscious minds. But why doesn’t all information processing go in the dark?
This question is the subject of long-standing debates across disciplines [seth2022]. One highly influential view is that consciousness has two aspects [seth2022]. The first is functional, by which we mean the ability to access and communicate information [block1995]. How information processing is linked to consciousness is the “easy problem” of consciousness [chalmers1995]. The second aspect is “what it is like” to consciously experience information processing, or phenomenal consciousness [chalmers1995, block1995, nagel1974, gallagher2021, fuchs2017]. This doesn’t mean just global states like being awake, but more specific local states like smelling a cup of coffee. These local contents or “qualia” are characterised by what it is like to be in them [seth2022]. It is unclear the extent to which functional and phenomenal aspects are independent. David Chalmers has influentially suggested that it may be possible to construct a “philosophical zombie” which acts in every way like a person but has no qualia [block1995]. For example, a thermostat certainly detects heat and so processes information, but there presumably is not anything it is like to be a thermostat. Hence the question “why is anything conscious” may be understood as “why is there sometimes a qualitative aspect to information processing?”. This is the “hard problem” of consciousness.
The hard problem has sparked a substantial body of work, detailed discussion of which [seth2022, northoff2014] lies beyond the scope of our paper. Rather, in what follows we build upon the useful distinction between lower and higher order theories of information processing in relation to conscious experiences.
For example, higher order thought theory (HOT) [rosenthal2005, brown2019] holds that the information of which a conscious being is aware are higher order “meta-representations” of lower order “local” mental states. Lower order states may include emotions and perceptions, while higher order meta-representations reflect upon those. The link between the two may explain something of the phenomenal character of states. Sense data is processed by the body resulting in lower order mental states, and then meta representations of those is where we might find more abstract conceptual or thought-like contents of consciousness.
The distinction between lower and higher orders is a good starting point to understand why most biological information processing goes on “in the dark”. Yet, the interesting question in our view is why do these lower order states arise? Can they occur in the absence of subjective, qualitative experience? It is widely agreed that higher order mental states such as desire to drink a coffee is accompanied by a qualitative aspect. But is there something it is like to “be in the dark”? That is: to process information at the lower bodily levels?
In this paper we suggest a way to dissolve the “hard problem” of consciousness is to reverse the order and start with the ‘impure’ embodied biological organism instead of the ‘pure’ abstract mental states.
If an explanation of how consciousness functions is to reveal why information processing sometimes has a qualitative aspect, then it must explain how we get local states [thompson2016], not begin by assuming local states. We need to go a level down and begin at the level of the embodied organism. Longtime considered a fringe approach, the embodied cognition paradigm [varela2016] has recently gained substantial influence in cognitive science and philosophy [ciaunica2023, levin2024, seth2018]. The idea is that instead of considering the body as a mere device designed to fuel and contain the mind (a device that can be replaced with a vat or a robot, for example), one must consider the mind as serving the self-sustaining needs of a surviving body.
If this is so then understanding consciousness must start with understanding the ‘humble’ lower bodily levels of information processing, not the higher order levels of information processing only. In other words, conscious experiences do not merely depend on bodily experiences as an external factor that can be replaced with a vat or an artificial system.
One promising approach to this question is to try to rigorously define consciousness from first principles and show that some aspects of functional consciousness depend on phenomenal consciousness in a manner that makes zombies impossible, “dissolving” the hard problem by showing the phenomenal to be intrinsically functional [^2] Note that we are far from the first to claim that the phenomenal is functional. For example, the proposed Conscious Turing Machine [blum2020] based on Global Workspace Theory [baars1997], and constructivist approaches to artificial intelligence [wang2020] take similar position. It is our explanation of how and why that is novel.. Unifying them. We would need to establish axioms that hold in every conceivable environment and show that it is impossible to have the function of consciousness without the subjective experience of it. Such an answer must explain how consciousness functions, and why most information processing goes on “in the dark”.
This paper lays the grounds of a computational model formalizing the link between lower and higher levels of information processing in relation to conscious experiences in a way that is compatible with the basic principles of embodied cognition and enactivism. Enactivism is roughly the view that information processing arises through a dynamic interaction between an acting organism and its environment [thompson2007].
Now, the notion of computation is widely debated, and a detailed review of these discussions would lead us to a major digression [piccinini2015, piccinini2021]. Here we define computation not in terms of abstract symbol shuffling and representation, but in concrete mechanistic terms where ‘information processing’ refers to the actual physical transition of a system from one state to another; a system such as a human body.
Pancomputationalism is the idea that all dynamic systems are constituted by computation, which for our purposes is defined as above. This paper builds upon the formalism of pancomputational enactivism [bennett2024a]. To answer “why is anything conscious” we proceed in two steps. First, we develop a formalism in which lower and higher order theories of consciousness, and phenomenal and access consciousness, are all derived from first principles. Each follows because of scaling natural selection pressures and the ability to adapt [^3] This is loosely inspired by a scale-based framing of machine learning [sutton2019] - we are applying Sutton’s “bitter lesson” to biology.. Second, we provide an argument building on that to dissolve the hard problem of consciousness, answering why there must a qualitative aspect to information processing and in what circumstances.
This paper is organised as follows. It should be noted first that the appendix at the end of this paper is a list of formal mathematical definitions, which the reader may refer to as needed. For readers without formal background, the appendix can be skipped. In Section 2 we extend separately published arguments, experiments and proof results [bennett2024a, bennett2023b, bennett2023c], developing a model of self-organising systems that holds across all conceivable environments. In Sections 3 and 4 we explain how this model formalises relevance realisation and unifies lower and higher order theories. Rather than assuming abstract objects are primary and trying to learn causal relations between them, this assumes valence is primary and that abstract objects are constructed to classify cause and anticipate valence. This takes us from attraction and repulsion to contentless states, through lower order thoughts to higher order meta representations. We call this the psychophysical principle of causality. In Section 5 we extend previous work [bennett2023c] on causal learning and the development of first (1ST), second (2ND) and third (3RD) order selves, and explain how they are constructed as a direct consequence of scaling the ability of a self-organising system to adapt with natural selection pressures. In Sections 6 and 7 we explain how subjective experience requires a 1ST order self, and conversely why a 1ST order self implies there is something it is like to be an organism that has a 1ST order self. We then argue that access consciousness requires the ability to communicate meaning, which requires both 1ST and 2ND order selves, and show how a philosophical zombie is impossible. In Section 8 we describe the development of consciousness as we scale up the capacity to adapt with natural selection pressures. We discuss examples of each stage from nematodes to humans. Section 9 is concluding remarks.
2 Back to Foundations
Rather than presupposing local states or an abstraction layer, we start from first principles to formalise all conceivable environments [bennett2024a, bennett2024b], as in definition 1. This is based on two assumptions:
Axiom 1: When there are things, we call these things the environment.
Axiom 2: Where things differ, we have different states of the environment.
Hence, the environment is a set $\Phi$ of contentless global states $\Phi$. We define a declarative “program” as a set of states and treat the states themselves as irreducible [^4] We do not concern ourselves with imperative programs, because there is an equivalence between declarative and imperative [howard1980].. The set $P$ contains all such programs. The powerset of $P$ is every aspect of the environment, because every aspect of any conceivable environment must be set of such programs, which is a subset of $P$.
Programs return true or false, so a program is a “fact” if it is true, and the current state of the environment is the set of all “facts”. Truth is determined with respect to states, and states are assumed to be mutually exclusive. For example, if time is one way in which things differ, then there is one state at a time. It follows that every aspect of the environment is a set of programs, and those programs are facts relative to any physical state they contain. An aspect of the environment is realized by a state if the programs it contains are facts.
When we say a state is contentless, we mean we assume nothing about what states contain or are. Their only significance is in how they relate to one another, and each such relation is a program. This is intentionally like to post-structuralism [eacho2020, hayles2005], in which the only meaning words have is given via their relations to each other. This is why previous work claims it applies to all conceivable environments [bennett2024a].
Universality Claim: Axioms 1 and 2 hold for every conceivable environment.
Intuitively, one might understand the set of all programs (the powerset of all states) as the set of all relations between these otherwise meaningless states.
2.1 Natural Selection and Embodiment
We then assume a process of natural selection. This produces embodied organisms.
A set of facts can represent anything [^5] Some may object, pointing out that this ignores composition. However, the application of a function is a fact regardless of whether it takes another function as input., hence the body of an organism must be a set of programs. We call this set a vocabulary, and it is formalised in definition 2 (see Appendix).
The vocabulary can be seen as hardware and software operating together. Piccinini [piccinini2015, piccinini2021] identifies two sorts of computation: abstract and concrete. Abstract computations mean or represent whatever we say they do. In contrast concrete computations are physical. They are the action of physical systems over time. The programs we talk about here are concrete, physical computations in that they are the relations between physical states. It follows that the computations a body performs are also concrete, and so no two organisms have the same vocabulary, because that would mean they are the same body.
This might seem inconvenient, as intuitively we might prefer to think about bodies as an abstract ‘type’. However, doing so would undermine any claim we might make, as earlier results in artificial intelligence (AI) show [bennett2023c, bennett2024c]. The meaning of abstractions depends upon interpretation, much like the behavior of software is determined by the abstraction layer that “interprets” it. At its least abstract software is machine code [^6] A “word” of machine code triggers a mechanistic process hardwired in the CPU by the human design. For example a word may copy the 32 bit value stored at memory address $X$ into the 32 bit register $Y$, which the next line of machine code “adds” to register $Z$ by looping over each bit in an “adder” circuit.. Machine code is interpreted by hardware, and that hardware determines every aspect of what that machine code does. A word of machine code is a mechanical trigger that only “means” whatever we have designed the hardware to do when we input that word. That hardware is an abstraction layer in which the software exists (including “higher level” abstraction layers like the Python interpreter). In other words, what we call software is nothing more than the state of hardware [bennett2024a]. This has undermined all but the most subjective of claims regarding the behavior of theorised, software superintelligence [hutter2010, leike2015, leike2018]. It is a flaw in the very idea of intelligent software. The distinction between software mind and hardware embodiment is subsequently called computational dualism [bennett2024a], because it is reminiscent of how traditional cartesianism conceived of a mental substance distinct from physical.
Here we suggest that must avoid computational dualism. If we are to dissolve the hard problem by starting with the ‘impure’ embodied biological organism instead of the ‘pure’ abstract mental states, then we cannot presuppose organisms use particular abstractions. Instead we must explain why particular abstractions are formed, in terms of contentless states, as the product of interaction between an organism and its environment [thompson2007, varela2016, rolla2021]. This process is sometimes called “relevance realisation” [vervaeke2012, vervaeke2013a, vervaeke2013b].
Some have argued enactivism is incompatible with computation because the set of possible abstractions that must be searched for relevance realisation is intractable [jaeger2024]. For tractability, one must first isolate a “small world” of relevant information from the “big world” of all information. Part of the reason we can address this is because we formalise embodiment, in concrete terms from the level of irreducible physical states. Embodiment implies finite resource constraints, so we adopt one more axiom from [bennett2024b].
Axiom 3: All aspects of the environment are spatially extended [bennett2024b].
This means that a body can occupy only a finite number of states, which is formalised in definition 2. A body is a vocabulary $\mathfrak{v}\subset P$ with finitely many elements. $\mathfrak{v}$ implies a formal language $L_{\mathfrak{v}}$ of interaction between body and environment. A statement in this formal language is an aspect of the environment. A set of “programs” from our vocabulary. Statements have truth conditions with respect to environmental states, and every statement has an extension. The extension of a statement is the set of all statements which are supersets of the first statement (the set of all other statements by which the first statement is implied). Extension is important we can relate statements by their truth conditions, forming a lattice. Most importantly, we avoid the distinction between software and hardware and so avoid computational dualism.
Embodiment is dispositional. As natural selection evolves a vocabulary, it begins to isolate a “small world” relevant to survival [jaeger2024, bennett2024b, friston2023], from the intractable “big world” $P$ of all information.
2.2 Self-Organising Systems as Self and World Constraints
Both snowflakes and human bodies are self-organising systems, yet we regard only the latter as conscious. What separates the two? To tackle this question, we need to introduce the notion of self and world constraints.
Given that we have an embodied formal language, we can talk about computation with inputs and outputs by just treating everything as embodied statements embedded and enacted within the environment. Given an input $i$, the set of all possible outputs is the extension $E_{i}$ of that input. This is because if $i$ is realised by the environment, then the environment is constrained to only those states that realise $i$, which constrains what other statements can be realised. We can use this fact to talk about “policies” as embodied [^7] As a state of hardware rather than software, which is how the term policy is normally used in reinforcement learning. constraints on behaviour. A policy is a statement whose extension constrains outputs, like a causal intermediary in functionalism [putnam1967]. If behaviour is “motivated”, then statements have valence determined by natural selection and only a subset of the statements a body could make are “fit”. As a result, an organism will self-organise to express some statements, but not others. Though our environmental states are contentless and their only significance is in how they relate to one another, natural selection must make them attractive or repulsive to a self-organising system. This is formalised by the $\mathfrak{v}$-task as in definition 3888A $\mathfrak{v}$-task is called because a $\mathfrak{v}$-task because it exists in the context of a vocabulary $\mathfrak{v}$.. If $\mathfrak{v}$ is the vocabulary of a body, and $\mathfrak{v}$-task $\mu=\langle I_{\mu},O_{\mu}\rangle$ is fit behaviour, then $I_{\mu}$ is all the statements that body can express in which it is possible to remain fit, and $O_{\mu}$ is all the statements that body can express in which it remains fit. The extension $E_{I_{\mu}}$ of $I_{\mu}$ would be every output that it is possible to choose given the inputs $I_{\mu}$, but only a subset of those $O_{\mu}\subset E_{I_{\mu}}$ are correct outputs. In other words, this defines a self-organising behaviour as a constraint on outputs, given inputs [^9] Conversely, a system which considered all outputs to be correct would not be self-organising..
2.3 Inference
Every statement in $L_{\mathfrak{v}}$ implies a constraint, because there are only so many outputs that can be expressed by a body at the same time as any given statement. Assume we want to constrain the body to $O_{\mu}$ given $I_{\mu}$. A body could express a statement $\pi$ (meaning $\pi$ is true), and $\pi$ could then constrain the body to only correct outputs $O_{\mu}\subset E_{I_{\mu}}$ if $O_{\mu}=E_{I_{\mu}}\cap E_{\pi}$. We call a constraining statement a policy. A policy constrains outputs given inputs. A correct policy is one that constrains outputs to only correct outputs. For the sake of intuition, think of “correct” as “fit” according to natural selection [^10] This need not always be the case.. This is formally expressed in definition 4.
“I survive therefore my model is viable.” - Mark Solms [solms2021]
If an organism remains alive, then its history is an ostensive definition of “fit” self-organising behaviour. That history of interactions is just a set of statements in the organism’s embodied language, which can be expressed as a $\mathfrak{v}$-task $\mathfrak{h}$. Each input and output in $\mathfrak{h}$ is behaviour the organism has exhibited and everything that was involved in that behaviour (the interaction between organism and environment). The set of all fit behaviour would then be another task $\mu$, and if the organism remains fit then $\mathfrak{h}\sqsubset\mu$ [^11] One task is a child of the other, meaning the inputs and outputs are subsets of the parents’ inputs and outputs..
2.4 Learning
Learning can be understood as the process of approximating the program which caused data. Constructing a “world” model. Here, we just need to explain how an organism gets $\mu$ from $\mathfrak{h}$. We don’t need a world model, but a way to constrain outputs to just those that are fit. Like everything else, that constraint is once again a statement in the formal language $L_{\mathfrak{v}}$.
“The best model of the world is the world itself” - Rodney Brooks [dreyfus2007]
Those examples in $\mathfrak{h}$ could have been generated by any one of a set of policies $\Pi_{\mathfrak{h}}$. These are not representations of possible causal intermediaries in the machine functionalist sense, but just parts of the outputs from which the outputs could be derived given the inputs. Every $\mathfrak{v}$-task (and thus every self-organising system) implies such a set of policies, and not all of those policies will constrain outputs the same way given new inputs. Some policies will “generalise” to imply fit behaviour in unfamiliar circumstances, meaning correct outputs given a new set of inputs (a parent task of the organism’s history $\mathfrak{h}$). If those policies imply fit behaviour given new inputs, then the organism will remain fit. If an organism learns, then it is adapting to embody a policy that it hopes will generalise to fit behaviour $\mu$.
Earlier work [bennett2023b] showed that the optimal strategy to learn and adapt as fast as possible [^12] Meaning to converge on a fit policy given the smallest history possible, to realise fit behaviour. is to prefer “weaker” policies, meaning those with larger extensions. The policy which generated outputs $O$ given inputs $I$ is most efficiently identified by constructing a policy $\pi$ such that $\pi$ generates $O$ from $I$, and $\pi$ implies the weakest constraint that can be implied while still generating $O$ from $I$. This maximises adaptability, because it allows a self-organising system to construct fit policies from a shorter history [bennett2023c]. This formalised in definition 5 and illustrated by figure 1.
Figure 1: Illustration of how child tasks are examples of parent tasks.
3 Relevance Realisation Through Causal Learning
When an embodied organism learns a policy, it learns a relevant interpretation of that information, labelling it through behaviour. To efficiently adapt, an organism must correctly anticipate valence, which means correctly identifying what causes valence.
For convenience, we will now expand upon the concept of organism given in definition 6. All organisms must have preferences, because they will make some decisions and not others. Behavior implies policies, and every policy implies tasks, so we can define preferences as an binary relation over tasks.
Natural selection favours adaptability, hence we assume it prefers organisims that prefer weaker policies, generalising to remain fit. When a system optimises for weak policies, we call it weak policy optimisation (WPO). If a policy is fit, then it must correctly predict causes of valence. Hence by constructing fit policies, an organism divides the world into relevant objects and properties. A weaker policy implies all more specific versions of itself, meaning those that more tightly constrain outputs by having a smaller extension. Hence policy learning implies a lattice of policies that vary in weakness. Every fit policy would be a “causal identity” for something relevant, for example a specific object like “this coffee” or a weaker and more general concept like “all coffee”.
We also need to introduce the notion of interaction between organisms. Organisms interact when they “affect” one another (in the physical rather than psychological sense of the word). This is formally defined in 9.
Now, Artificial Intelligence (AI) and machine learning [russell2005ai, bishop2006, sutton2018, goertzel2014] are concerned with engineering adaptive agents. In that context, causality has become a mainstream topic of research. Causal learning is demonstrably necessary to thrive in an interactive setting [pearl2018, richens2024robust]. Where the causal graph is known in advance (for example if we were to be measuring the efficacy of medical interventions), this issue could be resolved using causal language (such as “Do Calculus” [pearl2009]) to represent an intervention from outside the system the graph describes. Such an intervention represent agency, in that an organism observing the environment draws conclusion about that environment, and then intervenes to change the environment. To illustrate this point, consider the following example from [bennett2023c].
Suppose an organism named Bob is attempting to learn and predict the environment. Now assume Bob has observed Alice wearing a raincoat only when it rains, and that Bob has observed rain only at those times when Alice has been observed wearing a raincoat. If we represent the raincoat observation with a binary variable $O\in\{true,false\}$ and the advent of rain with a variable $R\in\{true,false\}$, and if Bob is a Bayesian, then Bob’s observations will lead to the conclusion that $p(R=true\mid O=true)=1$. This means Bob believes that if Alice wears a raincoat, then it must be raining. Now lets permit Bob to interact with its environment. Assume Bob wants it to rain. Based on the belief $p(R=true\mid O=true)=1$, Bob may conclude that forcing $O=true$ by holding a gun to Alice’s head and demanding Alice wear a raincoat will cause it to rain. This is absurd. $O=true$ does not represent the event $q=$“I coerced Alice into wearing a raincoat”, but an entirely different event $v=$“Alice decided to put on a raincoat for the same reason I have observed Alice wearing a raincoat in the past”. To accurately represent the environment, we need to represent that $p(R=true\mid q)=p(R=true)\neq p(R=true\mid v)=1$, meaning both $q$ and $v$. To illustrate the problem visually, we started with the acyclic graph
$O$$R$
and our intervention disconnected rain from the choice of clothing:
$O$$R$
This can be resolved by introducing a “do” operator that we apply to a variable $O$ to obtain $do[O=o]$, to represent the fact that an agency from outside the system has intervened to assign a value to $O$, so that we can represent $p(R=true\mid do[O=true])=p(R=true)\neq p(R=true\mid O=true)=1$. Thus, the aforementioned $q$ is equivalent to $do[O=true]$, while $v$ is equivalent to $O=true$.
The Psychophysical Principle of Causality
We have explained why we need to account for causality. However, we have not explained how one might come to know all the objects involved (where did the variables come from?), or how they relate to one another causally (to form a graph).
There are two ways to learn causal graphs. One can either assume a set of variables and learn relations between them, or assume relations and learn the objects that fit them [bennett2023c]. Beginning as we are with contentless states, we do the latter. Rather than assuming objects are primary and trying to learn causal relations between them, this assumes valence is primary and that abstract objects are constructed according to what causes valence. We call this the psychophysical principle of causality. This is what will take us from attraction to and repulsion from contentless physical states, through lower order thoughts to higher order meta representations. We start with the complete absence of content [thompson2016], and once a system learns policies there are distinct “contents”. Policies serve to classify objects and properties. Death grounds meaning.
To develop this idea, we first show adding additional nodes to a causal graph is an adequate substitute for the “do” operator [dawid2002].
Assume we again have Bob who constructs a causal graph of the environment. Assume Alice exists in that environment. From Bob’s perspective, Alice is just a part of the environment represented by a variable $A$ in Bob’s causal graph.
$O$$A$$R$
Now assume Bob observes Alice taking an action that changes an aspect of the environment, represented by the variable $O$ (for example, Bob observes Alice putting on the raincoat). From the perspective of Bob, Alice’s action is the assignation of a value to a variable $A$, and to the variable $O$. There is no need to involve a do operator in this scenario because we can already represent that $p(R=true\mid A=x,O=true)=p(R=true)\neq p(R=true\mid A=y,O=true)=1$ (because Alice is part of the causal graph). This raises the question; if we do not require a $do$ operator to represent the actions of Alice, then why would we need it to represent an intervention by Bob? Couldn’t we represent the same information by introducing a new variable? The answer is yes [bennett2023c, dawid2002].
$O$$B$$A$$R$
Now we will show how it is possible to learn the relevant nodes to add, in the form of statements. The matter of which cause-and-effect relations are learned is determined by valence, and so the objects learned are statements classifying causes and valence.
4 Relevant Causal Identities
The formalism we employ began as an attempt to explain why some abstractions are formed, and not others [bennettmaruyama2022a]. The optimal choice of policy for efficient adaptation is the weakest because it is implied by my statements in an embodied language. That language is dispositional, shaped by natural selection, and so a weaker policy is likely to be a fitter policy. As we have already shown, correct identification of cause and effect is often necessary to make accurate predictions and survive. WPO will thus correctly identify cause and effect where relevant. This is not just a case of substituting the “do” operator, but of simplifying the entire environment into objects and properties relevant to the organism’s motivations.
We will now explain how relevant objects are classified, as “causal identities” [bennett2023c] and protosymbols [bennett2023d].
4.1 Causal Learning
In accord with [bennett2023d, def. 6] we assume a vocabulary $\mathfrak{v}_{\mathfrak{b}}$ belonging to an organism $\mathfrak{b}$ (Bob). A “cause” in the context of this formalisation is not a variable but a statement $l\in L_{\mathfrak{v}_{\mathfrak{b}}}$ in that formal language. The raincoat example would involve $obs,rain\in L_{\mathfrak{v}_{\mathfrak{b}}}$ such that:
[…]
$int$$obs$$rain$
In other words, so long as $obs\neq int$, the intervention can be differentiated from the passive observation. We formally define this in definition 10.
This being the case, any set $c\subseteq int-obs$ could be used to identify the party undertaking the intervention, which is why $c$ is referred to as a “causal identity”. It distinguishes the intervention $int$ from the passively observed effect $obs$, like reafference in living organisms. However, the above only considers one intervention. A weaker or more general causal identity would be one that is shared by more interventions. For example, we might have two different interventions $a_{1}$ and $a_{2}$, with the observed effects $c_{1}\subset a_{1}$ and $c_{2}\subset a_{2}$, and causal identities $a_{1}-c_{1}=i_{1}$ and $a_{2}-c_{2}=i_{2}$. If $i_{3}=i_{1}\cap i_{2}\neq\emptyset$, then $i_{3}$ is a causal identity that is present in two interventions.
Being a question of classification, we can express a causal identity in relation to a $\mathfrak{v}$-task, for which the causal identity is a policy (expanding upon [bennett2023c, def. 10]), giving us definition 11.
Example 1.
Suppose we have organisms $\mathfrak{a}$ (Alice) and $\mathfrak{b}$ (Bob), and that the inputs Alice has experienced so far are $I_{\mathfrak{h}_{<t_{\mathfrak{a}}}}$. These can be divided into those in which Bob affected Alice $S^{\mathfrak{b}}_{\mathfrak{a}}$ and those in which Bob did not $S^{\lnot\mathfrak{b}}_{\mathfrak{a}}=I_{\mathfrak{h}_{<t_{\mathfrak{a}}}}-S^{\mathfrak{b}}_{\mathfrak{a}}$. By affecting Alice, Bob has intervened in Alice’s experience. Alice can construct a causal identity $b$ for Bob corresponding to interventions $INT=S^{\mathfrak{b}}_{\mathfrak{a}}$ and observations $OBS=S^{\lnot\mathfrak{b}}_{\mathfrak{a}}$.
4.2 Ascribing Intent to Other Objects
The “do” operator is used to discern the difference between an event one has caused, and that same event passively observed. However, what is passive observation if not the result of an intervention by something other than one’s self? The distinction between “intervention” and not is misleading. Everything is an intervention. The question is not “is this an intervention” but “by whom was this intervention made?”.
Earlier, we arrived at the following graph in which Bob’s intervention was given by $int$.
$int$$obs$$rain$
However, what if a third person Larry puts the coat on Alice? Surely Bob can observe this, and so Bob’s observation of Larry’s intervention is $v\in L_{\mathfrak{v}_{\mathfrak{a}}}$ such that $obs\subset v$. To account for this, Bob can construct a causal graph as below (with $b.$ representing Bob and $l.$ representing Larry).
$b.\ int$$l.\ int$$obs$$rain$
Bob’s causal identity for himself $c_{b}\subset b.\ int-obs$ only represents the intervention by himself. However, now we can see that Bob must also construct a causal identity $c_{l}$ for Larry, where the $c_{l}\subset l.\ int-obs$. More generally, for an organism $\mathfrak{a}$ with an embodied language $L_{\mathfrak{v}_{\mathfrak{a}}}$ to construct a causal identity for an object $\mathfrak{b}$, it must first be the case that $\mathfrak{a}$ is affected by $\mathfrak{b}$ [bennett2023d], to satisfy the incentive precondition for causal identity. It depends on motive, or valence. Recall that to be affected is formally defined in definition 9.
Assume an organism $\mathfrak{a}$ is affected by $\mathfrak{b}$ given inputs $INT$, and not affected given inputs $OBS$. To then attribute the contents of $INT$ to one specific entity, there must be something in common between the members of $INT$ caused by $\mathfrak{b}$ that is not shared by any member of $OBS$ caused by something else (in other words it must be at least possible for $\mathfrak{a}$ to discern the existence of $\mathfrak{b}$). The contents of $INT$ are “interventions” by $\mathfrak{b}$ and by learning $c$, a corresponding causal identity, $\mathfrak{a}$ can discern the existence of $\mathfrak{b}$. This is not to say that $\mathfrak{b}$ decides (chooses an output) or is even alive. What is important is that $\mathfrak{b}$ affects $\mathfrak{a}$, making it possible to discern when these interventions are a consequence of $\mathfrak{b}$’s existence.
4.3 Preconditions
There are preconditions for the existence of a causal identity. First, the vocabulary $\mathfrak{v}_{\mathfrak{a}}$ of an organism $\mathfrak{a}$ must be large and expressive enough to ensure that observations are distinguishable from interventions, meaning we have a vocabulary of sufficient scale.
Second, there must be an incentive to construct it. Inference is only possible if some states are preferable to others. It is a value judgement. As Hume pointed out, one cannot derive what “ought” to be from a statement of what “is”. Natural selection provides a notion of what ought to be, by eliminating anything which ought not.
  1. 1.
    The scale precondition requires $\mathfrak{v}$ contain the causal identity.
  2. 2.
    The incentive precondition is that fitness demands the causal identity.
4.4 Realising Lower Order States And Higher Order Meta Representations
The weaker the causal identity, the more pervasive the “identity” in the sense of being part of more interventions. Learning is not just about constructing policies but joining them together into weaker, more abstract policies. By learning policies that correctly identify the causes of valence at different levels of abstraction, the organism engages in relevance realisation.
This is formalised in definition 6. An organism $\mathfrak{o}$ has a set of policies it has learned. Each and every policy implies a $\mathfrak{v}$-task. A $\mathfrak{v}$-task is a triadic relation between inputs, outputs and policies which resembles Peircean semiosis [sep-peirce-semiotics, bennett2023d] of sign, referent and interpretant. Hence, in definition 7 we formalise this as a “protosymbol” [^13] Proto because it is something more primitive than a symbol as conceived of by Peirce. system $\mathfrak{s}_{\mathfrak{o}}$ for the organism $\mathfrak{o}$. Interaction is defined in terms of choosing an output, which means there is inevitably a policy and a policy implies a $\mathfrak{v}$-task, so definition 6 includes a preference order over tasks, which is used in definition 8 to formalise the interpretation of inputs (in terms of constraints, rather than an algorithm). In particular, the organism must always act according to an interpretation, and some interpretations imply others. Related as they are in a lattice, protosymbols are analogous lower order states and higher order meta representations. Tasks exist in a “generational hierarchy”. They are not mutually exclusive. Higher level tasks are more general, and have fewer policies because only very weak policies could complete them. Some have framed consciousness as a problem of moving from unary, to dyadic, to triadic relations [goertzel2006]. Similarly, we have gone from unary states to dyadic programs, to triadic tasks and protosymbols.
Figure 2: Illustration of how the generational hierarchy of tasks is a lattice of protosymbols.
The question now is when does an organism realise a causal identity for when it causes valence? When does it construct a causal identity for itself?
5 Multi-Layered Self-Organisation
As the vocabulary and capacity for WPO scale, a greater variety of concepts can be learned. Progressively higher orders of ‘causal-identity’ for one’s self related information processing are constructed [bennett2023c]. This lets us frame the construction of embodied selves in developmental [ciaunica2019] and evolutionary terms.
5.1 The First Order Self
A first order self (1ST henceforth, see figure 3) is functionally equivalent to reafference, which has been observed in the mammalian mid-brain and insect central cortex [merker2007, barron2016]. Reafference lets an organism discern the consequences of its actions, and so a 1ST order self is necessary for accurate inference in an interactive setting in the same sense as Pearl’s ‘do’ operator [pearl2018].
A 1ST order self formally defined in definition 13. It serves as the locus of self related information processing and experience [bennett2023c], allowing the organism to plan complex interactions and maintain a consistent “self” that is part of the present interaction. But also, to anticipate future planned interactions (for example, an insect navigating its environment [barron2016]), and recollect past interactions (it is a subset of all those relevant “statements” in the formal language).
Figure 3: Visual intuition for a 1ST order self. The organism constructs a causal identity for itself and can relate that to observed events.
Natural selection prefers efficiency. The absence of a single, centralised causal identity can create inefficiencies [^14] The effect of insufficiently weak representations can be observed in large language models like GPT-4 [bennett2023d, gpt], where they fail to make consistent statements about an object because represent it as multiple objects.. Decentralised or asynchronous control might be advantageous in some circumstances (e.g. an ant colony), but not others (e.g. in an individual human body). An organism that has a distributed control system might have a “self” for each part of that system, and perhaps the co-ordination of parts might seem to suggest the existence of a centralised “self”. Redundancy might be useful in some circumstances, but there is also a cost - more data may be required to learn and adapt, because the optimal choice of policy for accurate generalisation is the “weakest” [bennett2023b].
If the scale and incentive preconditions are met for a causal identity for one’s self, then the organism must learn one. This might be partly hard-coded, but our focus is on inference by the organism itself.
5.2 The Second Order Selves
Survival may demand organism $\mathfrak{a}$ infer $\mathfrak{b}$’s prediction of $\mathfrak{a}$’s interventions (to see one’s self as if through another’s eyes [bennett2023d]). This is called a second order self (2ND henceforth). We argue that if access conscious contents are available for communication in the human sense, then they must be communicable in the Gricean sense [grice1957, grice1969]. Grice argued that communication is about the inference of intent. If person $\mathfrak{a}$ and $\mathfrak{b}$ are talking, then the meaning $m_{\mathfrak{a}}$ of what $\mathfrak{a}$ says is whatever $\mathfrak{a}$ intends $\mathfrak{b}$ understand. The meaning $m_{\mathfrak{b}}$ that $\mathfrak{b}$ understands is whatever $\mathfrak{b}$ thinks $\mathfrak{a}$ wants $\mathfrak{b}$ to think. $\mathfrak{b}$ has understood what $\mathfrak{a}$ means if $m_{\mathfrak{b}}$ approximates $m_{\mathfrak{a}}$. This can happen only if $\mathfrak{a}$ can predict with reasonable accuracy what $\mathfrak{b}$ thinks $\mathfrak{a}$ thinks, and $\mathfrak{b}$ can predict what $\mathfrak{a}$ thinks $\mathfrak{b}$ will think upon hearing an utterance. In other words, both $\mathfrak{a}$ and $\mathfrak{b}$ must have 2ND order selves that are good approximations. Yes, there are other aspects to communication.
However, here we are talking about consciousness. Access conscious contents are those available for reasoning and report. It follows [^15] If the definition of access consciousness is to be consistent with reasoning and report as exhibited by conscious humans. that access conscious contents must in principle be communicable in the sense Grice described.
As such, we argue contents available for communication can only be the contents of 2ND order selves, which means only an organism with 2ND order selves can be considered to have access consciousness. 2ND order selves also explain attention and self-awareness. An organism can have many 2ND order selves because they depend upon who or what the organism is interacting with, just as the availability of information depends on context.
Where a 1ST order self might allow one to observe a cat and form plans regarding causal interactions with the cat, a 2ND order self would allow one to be consciously aware of the cat for the purpose of reasoning and report. One can know of the cat, and that another organism knows of the cat, but a 2ND order self is insufficient to be aware that one is aware of the cat [^16] This is ‘meta-self-reflexive consciousness’ as some have described it [morin2006]..
Figure 4: Visual intuition for a 2ND order self. The organism constructs a causal identity for itself, and for another object, and the causal identity for the other object includes a prediction of the organism itself from that object’s perspective. This would occur, for example, if one was a predator trying to predict the movements of prey in response to one’s own actions. It is an extension of the 1ST order self.
More formally using the notation given in definition 15, assume $\mathfrak{a}$ and $\mathfrak{b}$ are organisms that evolved to accurately predict one another’s behaviour. Assume $\mathfrak{a}$ constructs a causal identity $c^{\mathfrak{b}}_{\mathfrak{a}}$ to predict $\mathfrak{b}$ given input $i_{\mathfrak{a}}\in I_{\mu_{\mathfrak{a}}}$, of which a second order self $c^{\mathfrak{ba}}_{\mathfrak{a}}$ is part. Likewise, $\mathfrak{b}$ constructs $c^{\mathfrak{a}}_{\mathfrak{b}}$ to predict $\mathfrak{a}$ given input $i_{\mathfrak{b}}\in I_{\mu_{\mathfrak{b}}}$, of which $c^{\mathfrak{ab}}_{\mathfrak{b}}$ is part. What is important here is that each organism’s intent is to some extent inferred by the other, and that fact inevitably changes the sorts of policies that will be “fit”. For example, second order self means each knows the other can anticipate manipulation, which means the optimal policy will often be to have rather than feign intent that aligns to some extent with the other party’s desires, to co-operate [alexander2022] [^17] Depending upon circumstances, for example organisms may co-operate in some circumstances but not others [bennett2023d], and transient relations, information asymmetry and other factors can make deceit a more attractive option.. Repeated interaction creates an iterated prisoner’s dilemma, incentivising co-operation and signals that both parties interpret similarly (the beginnings of language) [bennett2023d]. To communicate in Gricean terms, $\mathfrak{a}$ must intend to convey meaning $m_{\mathfrak{a}}$, and $\mathfrak{b}$ must recognise this intent. The incentive precondition explains why $\mathfrak{a}$ would form such intent (co-operation is often advantageous), while how may be understood as follows:
  • $c^{\mathfrak{ba}}_{\mathfrak{a}}$ lets $\mathfrak{a}$ predict what $\mathfrak{b}$ will come to believe when it observes $\mathfrak{a}$’s behaviour.
  • $c^{\mathfrak{ab}}_{\mathfrak{b}}$ then lets $\mathfrak{b}$ predict what $\mathfrak{a}$ intends that $\mathfrak{b}$ believe.
$\mathfrak{a}$ can use $c^{\mathfrak{ba}}_{\mathfrak{a}}$ to infer behaviour to which $\mathfrak{b}$ will ascribe the intent to communicate $m_{\mathfrak{a}}$, and $c^{\mathfrak{ab}}_{\mathfrak{b}}$ lets $\mathfrak{b}$ infer that this is what $\mathfrak{a}$ intends. The “utterance” Grice refers to is how $\mathfrak{a}$ affects $\mathfrak{b}$ in accord with earlier definitions. Put another way, $\mathfrak{a}$ encodes $m_{\mathfrak{a}}$ into its behaviour in a manner that $\mathfrak{b}$ can decode (their respective second order selves act as encoders and decoders). By encode and decode, we mean a loose approximation of $m_{\mathfrak{a}}$ is communicated. There are of course shortcuts, for example of $\mathfrak{a}$ and $\mathfrak{b}$ are of the same species then they likely have similar motives and experiences, and so the efficient thing for each to do would be to use its own intent as an approximation of what the other might think. However, that does not obviate the need for second order selves, it just makes them easier to realise.
5.3 The Third Order Selves
We can continue scaling the vocabulary and capacity for WPO indefinitely, however for the purposes of this paper we will stop at 3RD order selves, because this appears to be the highest level of human consciousness, known as Meta self-awareness [morin2006]. This is the awareness that one is self-aware. If self-awareness stems from 2ND order selves, then it follows that meta self-awareness requires 3RD. Formally a 3RD order self for $\mathfrak{a}$ reflecting off $c^{\mathfrak{ba}}_{\mathfrak{a}}$ lets $\mathfrak{b}$ is $c^{\mathfrak{baba}}_{\mathfrak{a}}$. It is $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a}$.
Figure 5: Visual intuition for a 3RD order self. The organism constructs a 2ND order self for its 2ND order self, and so becomes aware that it is aware.
6 The What and Why of Consciousness
Up to now we have developed the conceptual toolbox that we use to dissolve the hard problem. We started with the observation that self-organizing systems such as a human body processes information to maintain itself in the face of change both within and without. An organism does not spring into existence understanding number systems or objects, but constructs them through interaction with the environment. One can regard them as policies according to which the environment is interpreted. To learn how to interpret the environment an organism must differentiate between states, must react to change, and learn policies according to the valence associated with that change.
We will now argue that: i) there is something it is like to be an organism with a 1ST order self, ii) that subjective experience is the process of learning and enacting a hierarchy or causal identities, and iii) that access consciousness requires both 1ST and 2ND order selves.
Because we do not presuppose representations, we explain how they emerge from attraction to and repulsion from contentless physical states. This means the foundation of information processing is valence.
As we saw earlier, learned policy has valence, and it is a classifier of inputs. Those inputs are “information” in the mechanistic sense of the environment being in one state and not another. Hence, a policy is a classifier of information, but that information is not in a language, and it is not yet something labelled or quantified (until there is a policy that labels or quantifies).
Clearly, prelinguistic self-organising systems must classify and attach value and disvalue to states and anticipated states to prioritise and make decisions. To learn is to sense “something” and, motivated by valence, construct a policy classifying that “something”. Physical states are contentless. In the absence of language, the difference between any two states can only be qualitative. In other words we hold that quality is fundamentally valenced, i.e. the product of actively sensing and being physically attracted and repelled within a given environment.
The bold claim here is that information processing (i.e. exploring one’s body and environment) in self-organising systems such as the human body is necessarily qualitative. Note that this is different from experience in the sense usually defined in the literature.
Experience in the sense of experiential phenomenological content cannot exist without experience in the sense of embodied exploration of the body and world as we defined here (see also Ciaunica in prep). To put it provocatively quality precedes quantity, and quantity is nothing more than the interpretation of quality. Quality comes first and experiences should be regarded on a continuum rather than a switch on/ switch off phenomena. All living systems experience the world through their bodies and as such, there is something what it is like to experience the world in that basic way (even when one is asleep or a baby). One can access those phenomenal, experiential aspects at a higher level, true, but by accessing them, it doesn’t mean that one ‘constructs’ consciousness or one becomes a conscious being. One is already consciously experiencing the world before one can explicitly access one’s own experiences in 2ND order selves.
It follows that every policy learned in this way must classify a quality. Hence every such policy is a local state. A causal identity imbued with disposition by valence. There would be a policy for one’s act of smelling coffee. For perceiving one’s friend. There would be something different it is like to interact with a hostile version of that very same friend. The 1ST order self accompanies everything an organism does. It has a quality, so the 1ST order self is “what it is like” to be that organism. Put another way, Nagel’s question of “what it is like to be” a particular organism could be answered if one could somehow have that organism’s first order self [nagel1974]. An organism does not make a decision to interpret information, it just reacts. What we call subjective experience is the process of learning and enacting a lattice of causal identities, and there is something that “has” all of those subjective experiences once there is a 1ST order self that is part of them all. This is where phenomenal consciousness begins, with experience as learning and exploration.
One’s 2ND order selves would also have a certain qualitative character, and one’s 3RD order too.
There would, however, be a very clear delineation between conscious and not. The absence of a 1ST order self would mean there would be no one policy we might point at to say “this is what it is like” to be an organism. No policy linking all interventions together, and so no “self” that experiences making them. Hence a 1ST order self must precede all others.
Our arguments align with those of Merker, Barron and Klein, who have linked subjective experience to reafference [merker2005, merker2007, barron2016]. We agree that reafference is key, but we provide a very different explanation of why and how. Their work presents biological evidence for subjective experience in organisms with reafference. In contrast, we derive the 1ST order self from first principles and explain why and how it is a classifies “what it is like” to be a particular organism. The 1ST order self also happens to be equivalent to reafference, so we arrive at the same conclusion as Merker from very different, mathematical premises. Hence these are complementary positions.
7 Unifying Lower and Higher Order Theories of Consciousness
The psychophysical principle of causality takes us from attraction to and repulsion from contentless physical states, up to higher order meta representations. This produces different orders of self. We suggest that the full richness of human subjective experience depends upon these different orders of self, as does human behavior [^18] We additionally speculate that different orders of learned self may interact through hierarchical planning in a distributed system, perhaps something like a connective core [shanahan2012]. Arguably [boltuc2012], the full richness of human consciousness may be distinguished from phenomenal and access consciousness alone..
Conscious “access” as a human has it involves meaningful report in the sense described by Grice, which as we’ve established requires 2ND order selves. This is very different from the mere “access” to information. Hence, we hold that contents of which one is “access conscious”, in the sense that they are available for reasoning and report, can only be the contents of 2ND order selves. However, supposing we could somehow contrive a system with 2ND order selves but no 1ST, then 2ND order selves alone would still not allow one to reason about how one’s interventions might affect the contents of 2ND order selves to communicate meaning. For that, one requires a 1ST order self. Hence access consciousness requires both 1ST and 2ND order selves.
Hence our radical and provocative claim is that phenomenal consciousness without access consciousness is likely very common, but the reverse is impossible. For example, Block holds that phenomenally conscious content is phenomenal, whereas access conscious content is representational. In pancomputational enactivism there are only contentless states and the programs they imply. Any abstract “representational” content is just organised phenomenal content, clustered according to what causes valence. We don’t discard these two sorts of consciousness, but unify them by showing how phenomenal consciousness gives rise to access.
Our framework thus makes a philosophical zombie impossible because there can be no perfect unconscious replica of a conscious human. The hard problem has it backwards. The question is not why qualia exist, but why anyone thinks representational contents can exist without first being learned through qualitative experience and discrimination.
Far from suggesting there is no such thing as qualia, this suggests instead that there is no such thing as purely representational content in anything but the tools we construct and our interpretations of their behaviour. Cells are a material with agency [ball2023]. A human is a “multiscale competency architecture” [fields2020, levin2021, levin2024]. Our intelligence is a swarm intelligence; the high level goal directed behaviour of a swarm of cells. When we embody human abstractions (e.g. arithmetic) in silico (e.g. the x86 instruction set), we disconnect the high level goal directed behaviour from the low level behaviour (and the motivating valence) that gave rise to it [bennett2024c]. We call the information embodied in the computer “representational” because it means something to us. However when we embody our abstractions in silico we disconnect them from the valence that motivated their construction [bennett2024b]. This suggests there is no such thing as representational contents. They are a fiction we have invented because we struggle to reduce our own abstractions to their basic nature: the causes of valence. This ‘dissolves’ the hard problem by going a level down to the fundamental drive: stay alive!
8 From Rocks to Einstein: The Hierarchy of Being
To illustrate how our argument applies in the real world we describe stages of conscious organism. Each stage follows from scaling up supply and demand for WPO, through natural selection (see figure 6). For each stage we point out animals which are likely to be at least so conscious:
$0:$ Unconscious (e.g. a rock)
$1:$ Hard Coded (e.g. protozoan)
$2:$ Learning (e.g. nematode)
$3:$ 1ST Order Self (e.g. housefly)
$4:$ 2ND Order Selves (e.g. cat)
$5:$ 3RD Order Selves (e.g. human)
Stage 0: Unconsciousness
Stage zero may be understood as the consciousness we are willing to assign to a rock: more exactly, the lack of consciousness so assigned.
  • Example: A rock.
Stage 1: Hard Coded
Stage one refers to adaptations hard-wired by natural selection. This is the sort of preset behaviour that allows complexity to persist [heylighen2023b] in a reasonably consistent environment.
  • What: Hard-coded adaptations. Habituation and sensitization.
  • How: The extension of fit behaviour is learned by natural selection and hard-coded into the organism as a policy (in DNA, form, the local environment etc).
  • Why: If the environment is very predictable, it may be more efficient to hard-code fit behavior.
  • Example: Single-celled protozoan.
Stage 2: Learning
Stage two introduces learning. To learn an organism must store, classify and order historical examples by valence. However there is not something it is like to be stage two, because there is no locus of “self”. A biological example of such a decentralised nervous system is the cubozoan box jellyfish Tripedalia cystophora. Even Tripedalia cystophora was recently shown to be capable of associative learning [bielecki2023]. An entirely distributed control system can “learn”. Likewise, stage two is exemplified by the nematode C. elegans [rankin2022, willett2018], which has a centralised nervous system and exhibits some ability to adapt with experience. However, the absence of a “self” prohibits cause and effect reasoning, which as others have already pointed out must limit spatial, navigational abilities [barron2016]. When starved C. elegans exhibit “increased locomotion and dispersal in a random, rather than directed, search” [barron2016, luersen2014, artyukhin2015], whereas something like a bee or an ant can recall and navigate to previously discovered food [wehner2013, seeley1995page295, oades1978]. A nematode can learn, but the absence of a self for the purpose of causal reasoning limits adaptability.
  • What: Learning.
  • How: Valence.
  • Why: An organism that can learn can survive in more circumstances than one that cannot.
  • Examples: Jellyfish, nematode.
Stage 3: 1ST Order Self
This is where phenomenal consciousness begins, with a 1ST order self. In biological terms this also implies reafference, which others have argued is the key to subjective experience, albeit for different reasons than what we have argued here [merker2005, merker2007, barron2016]. They identified a housefly as a good example of where subjective experience may begin, and we concur. We also hold this is where an organism might be said to have intent. Intuitively, the policy that motivated behaviour is the intent of that behaviour [^19] In the same way declarative and imperative programs are equivalent [howard1980].. The “weaker” a policy is, the more behaviours it motivates. Conversely, if one’s actions share anything in common it is the intent that motivated them. The more diverse the actions, the more “general” or “high level” the intent they share. For example, the action of eating tends to involve the intent of satisfying hunger. If a policy is implied by all of an organism’s behaviour, then it might have motivated that behaviour. While this is an unusual form of words, we contend that an organism is such a policy.
A stage three there is a self to feel simple hunger and pain. However, communication as described by Grice would be impossible [grice1969], because a stage three organism has no self-awareness, to represent another’s perception of their intent [bennett2023d]. Nor would a stage three organism be able to conceive of its own death, or shame, because it cannot conceive of itself.
  • What: A 1ST order self. Reafference. Phenomenal consciousness.
  • How: Embodiment in which intervention is not identical to observation.
  • Why: Accurate prediction of consequences of interventions. For example, a fly must distinguish between having moved, and the environment having moved, to navigate.
  • Example: Housefly.
Stage 4: 2ND Order Selves
Stage four is the 2ND order self, and importantly this is where we hold access consciousness begins because it is where information is available for report in the Gricean sense. In other words, access consciousness follows phenomenal consciousness. The ability of ravens to intentionally deceive [bugnyar2002] suggests they are at least stage four. Raven $\mathfrak{a}$, aware that it is being observed by raven $\mathfrak{b}$, will act as if it is hiding food in one location to mislead $\mathfrak{b}$, but will then move the food in another location unobserved by $\mathfrak{b}$. $\mathfrak{a}$ seems to predict not just the intent of $\mathfrak{b}$ (to steal the food), but $\mathfrak{b}$’s perception of $\mathfrak{a}$. It seems intuitively likely that dogs and cats have second order selves, as they must hunt reasonably intelligent animals and must anticipate how their actions are perceived. For example, a cat anticipates its prey will flee when it is observed, and hides.
  • What: Access consciousness. Theory of mind. Self-awareness. Inner narrative.
  • How: Selection pressures that demand theory of mind.
  • Why: A 2ND order self is necessary to anticipate, manipulate and communicate intent. It allows one to anticipate how others will perceive one’s behaviour. This could be anticipating the reactions of predator or prey, or navigating a social hierarchy.
  • Example: Cats, dogs, ravens.
Stage 5: 3RD Order Selves
A 3RD order self facilitates complex planning, particularly of a social nature. A 3RD order self is a 2ND order self for one’s 2ND order self. Humans appear to possess this, because we are aware that we are aware. The seemingly altruistic behavior observed in Australian magpies [potvin2022] suggests they may also have 3RD order selves.
  • What: Meta self-awareness. Inner narrative in which actors have inner narratives.
  • How: More accurate prediction and planning.
  • Why: Because a social organism must predict complex social dynamics.
  • Example: Human. Highly intelligent animals such as Australian magpies may also be this conscious.
Figure 6: Overview of stages and orders of self.
9 Conclusions: Why Nature Does Not Like Zombies
In this paper we formally unified lower order states and higher order meta-representations, dissolving the hard problem of consciousness. We described a multilayered formalism illustrating how biological self-organising systems become phenomenally conscious when they construct a 1ST order self. A human lacking a 1ST order self could not perform causal reasoning needed to adapt as humans evidently can [pearl2018].
As previous research pointed out [bennett2024a], the computer metaphor with the seminal distinction between hardware and software is a simplification. Rather, software is a state of hardware. Nature privileges efficiency over abstract simplification. Compare the vast quantities of both training data and energy required by a large language model, to the small quantities humans need to solve a problem. Biological systems are more efficient, because they are adaptive at a every level [levin2024, bennett2024c]. Hence, rather than trying to explain the mind in the abstract like software, we started at the level of the embodied organism.
One consequence of our approach is that it places the phenomenal quality of conscious experience before access consciousness. We have shown a consistent definition of access consciousness requires 1ST and 2ND order selves. Phenomenal consciousness arises first, access comes later. Unlike panpsychism, we don’t believe rocks are conscious. Only self-organising systems that need to adapt, motivated by valence, while keeping track of the self are conscious. Consciousness is a necessary adaptation, and a philosophical zombie is impossible.
Furthermore, the conceivability of zombies depends on the ability to duplicate disposition but not phenomenal experience. Carruth argued that if qualia are dispositional, then that is impossible [carruth2016]. Our arguments align with Carruth’s. A causal identity is inherently dispositional, as is embodiment.
In summary, we have argued that access consciousness at the human level is impossible without the ability to hierarchically model i) the self, ii) the world and others and iii) the self as modelled by others. These follow from learning and enacting a lattice of causal identities, if the scale and incentive preconditions are met. There is something it is like to be an organism with a self, and that is where phenomenal consciousness begins. Our proposal lays the foundation of a formal science of consciousness, deeply connected with natural selection rather than abstract thinking, closer to the organic facts of humanity than the conceptual fiction of zombies.
Compliance with Ethical Standards
Funding
This work was supported by a Fundaçao para a Ciencia e a Tecnologia (FCT) grant PTDC/FER-FIL/4802/2020 ; and 2020-02773 CEECIND FCT to A.C.
Ethics Approval
This article does not contain any studies with human participants or animals performed by any of the authors.
Appendix
For convenience of reference, we have placed all definitions here. We have aimed to make the gist of the paper understandable without the math, hence the reader can either skip this section or refer back to it if needed. Many of the definitions have been adapted from a variety of preceding work [bennett2024a, bennett2023b, bennett2023c, bennett2023d, bennett2024b, bennett2024c]. They are referred to in the body of the paper when they become relevant, in the order in which they appear here.
Definition 1 (environment).
  • We assume a set $\Phi$ whose elements we call states.
  • A declarative program is $f\subseteq\Phi$, and we write $P$ for the set of all programs (the powerset of $\Phi$).
  • By a truth or fact about a state $\phi$, we mean $f\in P$ such that $\phi\in f$.
  • By an aspect of a state $\phi$ we mean a set $l$ of facts about $\phi$ s.t. $\phi\in\bigcap l$. By an aspect of the environment we mean an aspect $l$ of any state, s.t. $\bigcap l\neq\emptyset$. We say an aspect of the environment is realised [^20] Realised meaning it is made real, or brought into existence. by state $\phi$ if it is an aspect of $\phi$.
Definition 2 (abstraction layer).
  • We single out a subset $\mathfrak{v}\subseteq P$ which we call the vocabulary of an abstraction layer. The vocabulary is finite.
  • $L_{\mathfrak{v}}=\{l\subseteq\mathfrak{v}:\bigcap l\neq\emptyset\}$ is a set of aspects in $\mathfrak{v}$. We call $L_{\mathfrak{v}}$ a formal language, and $l\in L_{\mathfrak{v}}$ a statement.
  • We say a statement is true given a state iff it is an aspect realised by that state.
  • A completion of a statement $x$ is a statement $y$ which is a superset of $x$. If $y$ is true, then $x$ is true.
  • The extension of a statement $x\in{L_{\mathfrak{v}}}$ is $E_{x}=\{y\in{L_{\mathfrak{v}}}:x\subseteq y\}$. $E_{x}$ is the set of all completions of $x$.
  • The extension of a set of statements $X\subseteq{L_{\mathfrak{v}}}$ is $E_{X}=\bigcup\limits_{x\in X}E_{x}$.
  • We say $x$ and $y$ are equivalent iff $E_{x}=E_{y}$.
(notation) $E$ with a subscript is the extension of the subscript212121e.g. $E_{l}$ is the extension of $l$..
(intuitive summary) $L_{\mathfrak{v}}$ is everything which can be realised in this abstraction layer. The extension $E_{x}$ of a statement $x$ is the set of all statements whose existence implies $x$, and so it is like a truth table. Intuitively a sensorimotor system is an abstraction layer. Likewise, a computer (each statement asserting a state of the computer, or a part thereof).
Definition 3 ($\mathfrak{v}$-task).
For a chosen $\mathfrak{v}$, a task $\alpha$ is a pair $\langle{I}_{\alpha},{O}_{\alpha}\rangle$ where:
  • ${I}_{\alpha}\subset L_{\mathfrak{v}}$ is a set whose elements we call inputs of $\alpha$.
  • ${O_{\alpha}}\subset E_{I_{\alpha}}$ is a set whose elements we call correct outputs of $\alpha$.
${I_{\alpha}}$ has the extension $E_{I_{\alpha}}$ we call outputs, and ${O_{\alpha}}$ are outputs deemed correct. $\Gamma_{\mathfrak{v}}$ is the set of all tasks given $\mathfrak{v}$.
[…]
(notation) If $\omega\in\Gamma_{\mathfrak{v}}$, then we will use subscript $\omega$ to signify parts of $\omega$, meaning one should assume $\omega=\langle{I}_{\omega},{O}_{\omega}\rangle$ even if that isn’t written.
(intuitive summary) To reiterate and summarise the above:
  • An input is a possibly incomplete description of a world.
  • An output is a completion of an input [def. 2].
  • A correct output is a correct completion of an input.
A $\mathfrak{v}$-task is a formal, behavioural description of goal directed behaviour. For example, an organism could be described by all behaviour in which it remains alive. Likewise, a $\mathfrak{v}$-task could describe a Turing machine.
Definition 4 (inference).
  • A $\mathfrak{v}$-task policy is a statement $\pi\in L_{\mathfrak{v}}$. It constrains how we complete inputs.
  • $\pi$ is a correct policy iff the correct outputs $O_{\alpha}$ of $\alpha$ are exactly the completions $\pi^{\prime}$ of $\pi$ such that $\pi^{\prime}$ is also a completion of an input.
  • The set of all correct policies for a task $\alpha$ is denoted $\Pi_{\alpha}$. [^22] To repeat the above definition in set builder notation: ${\Pi_{\alpha}}=\{\pi\in L_{\mathfrak{v}}:{E}_{I_{\alpha}}\cap E_{\pi}={O_{\alpha}}\}$
Assume $\mathfrak{v}$-task $\omega$ and a policy $\pi\in L_{\mathfrak{v}}$. Inference proceeds as follows:
  1. 1.
    we are presented with an input ${i}\in{I}_{\omega}$, and
  2. 2.
    we must select an output $e\in E_{i}\cap E_{\pi}$.
  3. 3.
    If $e\in{O}_{\omega}$, then $e$ is correct and the task “complete”. $\pi\in{\Pi}_{\omega}$ implies $e\in{O}_{\omega}$, but $e\in{O}_{\omega}$ doesn’t imply $\pi\in{\Pi}_{\omega}$ (an incorrect policy can imply a correct output).
(intuitive summary) To reiterate and summarise the above:
  • A policy constrains how we complete inputs.
  • A correct policy is one that constrains us to correct outputs.
In functionalist terms, a policy is a “causal intermediary” between inputs and outputs.
Definition 5 (learning).
. A proxy $<$ is a binary relation on statements. In this paper we use only one proxy, called the weakness proxy, which compares the cardinality of a statement’s extension. For statements $l_{1},l_{2}$ we have $l_{1}<l_{2}$ iff $\lvert Z_{l_{1}}\rvert<\lvert Z_{l_{2}}\rvert$. Whenever we use $<$ to compare statements, we are referring to the aforementioned weakness proxy.
(generalisation) A statement $l$ generalises to a $\mathfrak{v}$-task $\alpha$ iff $l\in\Pi_{\alpha}$. We speak of learning $\omega$ from $\alpha$ iff, given a proxy $<$, $\pi\in{\Pi}_{\alpha}$ maximises $<$ relative to all other policies in ${\Pi}_{\alpha}$, and $\pi\in{\Pi}_{\omega}$.
(probability of generalisation) We assume a uniform distribution over $\Gamma_{\mathfrak{v}}$. If $l_{1}$ and $l_{2}$ are policies, we say it is less probable that $l_{1}$ generalizes than that $l_{2}$ generalizes, written $l_{1}<_{g}l_{2}$, iff, when a task $\alpha$ is chosen at random from $\Gamma_{\mathfrak{v}}$ (using a uniform distribution) then the probability that $l_{1}$ generalizes to $\alpha$ is less than the probability that $l_{2}$ generalizes to $\alpha$.
(sample efficiency) Suppose $\mathfrak{app}$ is the set of all pairs of policies. Assume a proxy $<$ returns $1$ iff true, else 00. Proxy $<_{a}$ is more sample efficient than $<_{b}$ iff
$\left(\sum_{(l_{1},l_{2})\in\mathfrak{app}}\lvert(l_{1}<_{g}l_{2})-(l_{1}<_{a}l_{2})\rvert-\lvert(l_{1}<_{g}l_{2})-(l_{1}<_{b}l_{2})\rvert\right)<0$
(optimal proxy) There is no proxy more sample efficient than weakness. The weakness proxy formalises the idea that “explanations should be no more specific than necessary” (see Bennett’s razor in [bennett2023b]).
(intuitive summary) Learning is an activity undertaken by some manner of intelligent agent, and a task has been “learned” by an agent that knows a correct policy. Humans typically learn from “examples”. An example of a task is a correct output and input. A collection of examples is a child task, so “learning” is an attempt to generalise from a child to one of its parents. The lower level the child from which an agent generalises to parent, the “faster” it learns, the more sample efficient the proxy. The most sample efficient proxy is weakness [bennett2023b, prop. 1, 2], which is why we’re using it here.
Definition 6 (organism).
We can describe the circumstances of an organism $\mathfrak{o}$ as $\langle\mathfrak{v}_{\mathfrak{o}},\mu_{\mathfrak{o}},\mathfrak{p}_{\mathfrak{o}},<_{\mathfrak{o}}\rangle$ where:
  • ${O}_{\mu_{\mathfrak{o}}}$ contains every output which qualifies as “fit” according to natural selection.
  • $\mathfrak{p}_{\mathfrak{o}}$ is the set of policies an organism knows, s.t. $\mathfrak{p}_{\mathfrak{o}}\subset\mathfrak{p}_{n.s.}\cup\mathfrak{p}_{\mathfrak{h}_{{<t}_{\mathfrak{o}}}}$ and:
    • $\mathfrak{p}_{n.s.}\subset L_{\mathfrak{v}_{\mathfrak{o}}}$ is reflexes hard coded from birth by natural selection.
    • $\mathfrak{p}_{\mathfrak{h}_{{<t}_{\mathfrak{o}}}}=\underset{\zeta\in\mathfrak{h}_{{<t}_{\mathfrak{o}}}}{\bigcup}\Pi_{\zeta}$ is the set of policies it is possible to learn from a history of past interactions represented by a task $\mathfrak{h}_{{<t}_{\mathfrak{o}}}$.
    • If $\Pi_{{\mathfrak{h}_{<t_{\mathfrak{o}}}}}\not\subset(\mathfrak{p}_{\mathfrak{o}}-\mathfrak{p}_{n.s.})$ then the organism has selective memory. It can “forget” outputs, possibly to productive ends if they contradict otherwise good policies.
  • $<_{\mathfrak{o}}$ is a binary relation over $\Gamma_{\mathfrak{v}_{\mathfrak{o}}}$ we call preferences.
(intuitive summary) Strictly speaking an organism $\mathfrak{o}$ would be a policy, but we can describe the circumstances of its existence as a task $\mu$ that describes all “fit” behaviour for that organism. We can also identify policies the organism “knows”, because these are implied by the policy that is the organism. Likewise, we can represent lossy memory by having the organism “know” fewer policies than are implied by its history of interactions. Finally, preferences are the particular “protosymbol” the organism will use to “interpret” an input in later definitions.
Definition 7 (protosymbol system).
Assume an organism $\mathfrak{o}$. For each policy $p\in\mathfrak{p}_{\mathfrak{o}}$ there exists a set $\mathfrak{s}_{p}=\{\alpha\in\Gamma_{\mathfrak{v}_{\mathfrak{o}}}:p\in\Pi_{\alpha}\}$ of all tasks for which $p$ is a correct policy. The union of all such sets is
$\mathfrak{s}_{\mathfrak{o}}=\underset{p\in\mathfrak{p}_{\mathfrak{o}}}{\bigcup}\{\alpha\in\Gamma_{\mathfrak{v}_{\mathfrak{o}}}:p\in\Pi_{\alpha}\}$
We call $\mathfrak{s}_{\mathfrak{o}}$ a “protosymbol system”. A $\mathfrak{v}$-task $\alpha\in\mathfrak{s}_{\mathfrak{o}}$ is called a “protosymbol”, and is “more abstract” if it is higher in the generational hierarchy.
Definition 8 (interpretation).
Interpretation is an activity undertaken by an organism $\mathfrak{o}=\langle\mathfrak{v}_{\mathfrak{o}},\mu_{\mathfrak{o}},\mathfrak{p}_{\mathfrak{o}},<_{\mathfrak{o}}\rangle$, as follows:
  1. 1.
    Assume an input $i\in L_{\mathfrak{v}_{\mathfrak{o}}}$.
  2. 2.
    We say that $i$ signifies a protosymbol $\alpha\in\mathfrak{s}_{\mathfrak{o}}$ if $i\in I_{\alpha}$.
  3. 3.
    $\mathfrak{s}_{\mathfrak{o}}^{i}=\{\alpha\in\mathfrak{s}_{\mathfrak{o}}:i\in I_{\alpha}\}$ is the set of all protosymbols which $i$ signifies.
  4. 4.
    If $\mathfrak{s}_{\mathfrak{o}}^{i}\neq\emptyset$ then $i$ means something to the organism in the intuitive sense that there is “affect” or “value” compelling the organism to act.
  5. 5.
    If $i$ means something, then $\mathfrak{o}$ chooses $\alpha\in\mathfrak{s}_{\mathfrak{o}}^{i}$ that maximises its preferences $<_{\mathfrak{o}}$.
  6. 6.
    The organism then infers an output $o\in E_{i}\cap E_{\Pi_{\alpha}}$.
(intuitive summary) Interpretation is inference, with the additional step of choosing policies according to preference. This allows for irrational and instinctive choices, as well as rational ones. Intuitively, $i$ is every aspect of the context in which the organism finds itself; everything that can influence its interpretation.
Definition 9 (to affect).
Suppose we have two organisms, $\mathfrak{a}$ (Alice) and $\mathfrak{b}$ (Bob). Suppose $\mathfrak{a}$ interprets $i\in L_{\mathfrak{v}_{\mathfrak{o}}}$ as an output $o$, then:
  • a statement $v\subset i$ affects $\mathfrak{a}$ if $\mathfrak{a}$ would have interpreted $e=i-v$ as a different output $g\neq o$.
  • an organism $\mathfrak{b}$ has affected $\mathfrak{a}$ by making an output $k$ if, because of $k$, there exists $v\subset s$ which affects $\mathfrak{a}$.
Definition 10 (intervention).
By event we mean a statement in $L_{\mathfrak{v}}$, and an event happens or is observed iff it is a true statement given a state $\phi$. If $obs\in L_{\mathfrak{v}}$ is sensorimotor activity we interpret as an “observed event”, and $int\in L_{\mathfrak{v}}$ is an intervention to cause that event, then $obs\subset int$ (because $int$ could not be said to cause $obs$ unless $obs\subset int$).
(intuitive summary) An intervention is action undertake an organism or other agency, in the sense described by Pearl [pearl2018]. Intuitively, if “$int$” and “$obs$” are events which have happened, then we say that $int$ has caused $obs$ if $obs$ would not have happened in the absence of $int$ (counterfactual).
Definition 11 (causal identity).
If $obs\in L_{\mathfrak{v}}$ is an observed event, and $int\in L_{\mathfrak{v}}$ is in intervention causing $obs$, then $c\subseteq int-obs$ “identifies” or “names” the intervening agency. If $c=\emptyset$ then we have no way of knowing the intervening agency, if there is one. We call $c$ a causal identity corresponding to $int$ and $obs$. Suppose $INT$ and $OBS$ are sets of statements, and we assume $OBS$ contains observed events and $INT$ interventions, then a causal identity corresponding to $INT$ and $OBS$ is $c\neq\emptyset$ s.t. $\forall i\in INT(c\subset int)$ and $\forall obs\in OBS(c\cap obs=\emptyset)$ (we can attempt to construct a causal identity for any $INT$ and $OBS$). If a policy is a causal identity, then the associated task is to classify interventions.
Definition 12 (purpose, goal or intent).
We consider a policy $c$ which is a causal identity corresponding to $INT$ and $OBS$ to be the intent, purpose or goal ascribed to the interventions. $c$ is what the interventions share in common, meaning the “name” or “identity” of behaviour is the “intent”, “goal” or “purpose” of behaviour. Just as an intervention caused an observation, the intent which motivated the agency undertaking the intervention is what caused it (to correctly infer intent, one must infer a causal identity that implies subsequent interventions).
Definition 13 (1ST order self).
If $c$ is the lowest level causal identity corresponding to $INT$ and $OBS$, and $INT$ is every intervention an organism could make (not just past interventions, but all potential future interventions), then we consider $c$ to be the system’s 1ST order self. If $c\in\mathfrak{p}_{\mathfrak{o}}$ then an organism has constructed a 1ST order self. A 1ST order self for an organism $\mathfrak{o}$ is denoted $\mathfrak{o^{1}}$. An organism has at most one 1ST order self.
(intuitive summary) Intuitively, $\mathfrak{o^{1}}$ is where we draw the line between what the organism can intend and what it cannot. It is conceivable we might have two “organisms” in the same body by this definition, each with its own 1ST order causal identity. Ultimately, where an organism begins or ends remains malleable.
Definition 14 (preconditions).
If $\mathfrak{o}$ is an organism, and $c$ is a causal identity, the $\mathfrak{o}$ will construct $c$ only if the representation and incentive preconditions below are met:
  • the scale precondition is met iff $c\in L_{\mathfrak{v_{o}}}$, and
  • the incentive precondition is met if $\mathfrak{o}$ must learn $c$ to remain “fit”.
(intuitive summary) If $c$ is a 1ST order self, then these are the preconditions that must be met for an organism to construct $c$. Likewise, any other sort of causal identity.
Definition 15 (chain notation).
Suppose we have two organisms, $\mathfrak{a}$ (Alice) and $\mathfrak{b}$ (Bob). $c^{\mathfrak{b}}_{\mathfrak{a}}$ denotes a causal identity for $\mathfrak{b}$ constructed by $\mathfrak{a}$ (what Alice thinks Bob intends). Subscript denotes the organism who constructs the causal identity, while superscript denotes the object. The superscript can be extended to denote chains of predicted causal identity. For example, $c^{\mathfrak{ba}}_{\mathfrak{a}}\subset c^{\mathfrak{b}}_{\mathfrak{a}}$ denotes $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a^{1}}$ (what Alice thinks Bob thinks Alice intends). The superscript of $c^{\mathfrak{*}}_{\mathfrak{a}}$ can be extended indefinitely to indicate recursive predictions; however the extent recursion is possible is determined by $\mathfrak{a}$’s vocabulary $\mathfrak{v}_{\mathfrak{a}}$. Finally, Bob need not be an organism. Bob can be anything for which Alice constructs a causal identity.
[…]
An $n^{th}$ order self for $\mathfrak{a}$ is $\mathfrak{a^{n}}=c^{\mathfrak{*a}}_{\mathfrak{a}}$ where * is replaced by a chain, and $n$ denotes the number of reflections. For example, a 2ND order self $\mathfrak{a^{2}}=c^{\mathfrak{ba}}_{\mathfrak{a}}$, and a 3RD order self $\mathfrak{a^{3}}=c^{\mathfrak{baba}}_{\mathfrak{a}}$. We use $\mathfrak{a^{2}}$ to refer to any 2ND order self, and chain notation to refer to a specific 2ND order self, for example $c^{\mathfrak{ba}}_{\mathfrak{a}}$. The union of two $n^{th}$ order selves is also considered to be an $n^{th}$ order self, for example $\mathfrak{a^{3}}=c^{\mathfrak{baba}}_{\mathfrak{a}}\cup c^{\mathfrak{dada}}_{\mathfrak{a}}$, and the weaker or higher level a self is in the generational hierarchy, the more selves there are of which it is part.
Definition 17 (stages of consciousness).
We argue the following stages by scaling the ability to learn weak policies:
  1. 1.
    Hard Coded: organism that acts but does not learn, meaning $\mathfrak{p}_{\mathfrak{o}}$ is fixed from birth.
  2. 2.
    Learning: an organism that learns, but $\mathfrak{o^{1}}\not\in\mathfrak{p}_{\mathfrak{o}}$ either because $\mathfrak{o^{1}}\not\in L_{\mathfrak{v_{o}}}$ (failing the “scale precondition”) or because the organism is not incentivised to construct $\mathfrak{o^{1}}$ (failing the “incentive precondition”).
  3. 3.
    1ST order self: reafference and phenomenal or core consciousness are achieved when $\mathfrak{o^{1}}\in\mathfrak{p}_{\mathfrak{o}}$ is learned by an organism because of attraction to and repulsion from statements in $L_{\mathfrak{v}_{\mathfrak{o}}}$.
  4. 4.
    2ND order selves:
    1. (a)
      access or self-reflexive consciousness is achieved when $\mathfrak{o^{2}}\in\mathfrak{p}_{\mathfrak{o}}$.
    2. (b)
      hard consciousness is achieved when a phenomenally conscious organism learns a 2ND order self (an organism is consciously aware of the contents of 2ND order selves, which must have quality if learned through phenomenal conscious).
  5. 5.
    3RD and higher order selves: meta self-reflexive consciousness (human level hard consciousness) is achieved when $\mathfrak{o^{3}}\in\mathfrak{p}_{\mathfrak{o}}$.
Proposition
An organism that uses weakness as its proxy will learn an $n^{th}$ order self if the incentive and scale preconditions are met for that order of self.
Proof sketch. Assume we have an organism $\mathfrak{a}$ that learns using “weakness” as a proxy. A $\mathfrak{v}_{\mathfrak{a}}$-task $\mathfrak{h}_{<t_{\mathfrak{a}}}$ represents the history of $\mathfrak{a}$ (meaning $\mathfrak{h}_{<t_{\mathfrak{a}}}\sqsubset\mu_{\mathfrak{a}}$ and $\mathfrak{h}_{<t_{\mathfrak{a}}}$ is an ostensive definition of $\mu_{\mathfrak{a}}$, because $\mathfrak{a}$ remains alive). The organism explores the environment, intervening to maintain homeostasis. As it does so, more and more inputs and outputs are included in $\mathfrak{h}_{<t_{\mathfrak{a}}}$. It follows that:
  1. 1.
    From the scale precondition we have that there exists a $n^{th}$ order self $\mathfrak{a^{n}}\in L_{\mathfrak{v}_{\mathfrak{a}}}$.
  2. 2.
    To remain fit, $\mathfrak{a}$ must “generalise” to $\mu_{\mathfrak{a}}$ from $\mathfrak{h}_{<t_{\mathfrak{a}}}$. According to the incentive precondition, generalisation to $\mu_{\mathfrak{a}}$ requires $\mathfrak{a}$ learn the $n^{th}$ order self, which is when $\mathfrak{a^{n}}\in\mathfrak{p_{\mathfrak{a}}}$.
  3. 3.
    From [bennett2023b, prop. 3] we have proof that weakness is the optimal choice of proxy to maximise the probability of generalisation from child to parent is the weakest policy. It follows that $\mathfrak{a}$ will generalise from $\mathfrak{h}_{<t_{\mathfrak{a}}}$ to $\mu_{\mathfrak{a}}$ given the smallest history of interventions with which it is possible to do so (meaning the smallest possible ostensive definition, or cardinality $\lvert D_{\alpha}\rvert$).
Were we to assume learning under the above conditions does not construct an $n^{th}$ order self for $\mathfrak{a}$, then one of the three statements above would be false and we would have a contradiction. It follows that the proposition must be true. $\square$